OpenStudy (youngnephew34):

What is the geometric mean of 5 and 12? A) 60 B) 8 1/2 C) 2 square root of 15 D) 30 Someone HELP ME PLZZZZZ

OpenStudy (youngnephew34):

you there

OpenStudy (jack1):

use this example to calculate yours: mean of 4,11 and 14... arithmetic mean of 4 and 11 and 14 = add the terms... then divide by the number of terms (4 + 11 + 14)/ 3 terms = 29/3 geometric mean of 4 and 11 and 14 = multiply the terms... then find the nth root where n is the number of terms cube root of (4 x 11 x 14) = cube root of 616

OpenStudy (jack1):

note: cube root because there are 3 terms... cube root looks like this: \[\sqrt[3]{616}\]

OpenStudy (jack1):

so does that make sense now?

OpenStudy (youngnephew34):

yea Igot 24.8

OpenStudy (jack1):

ok... what about for yours tho?

OpenStudy (youngnephew34):

wym

OpenStudy (jack1):

geometric mean of 5 and 12 = sqrt (5 x 12) =?

OpenStudy (youngnephew34):

well 5x12=60 and the square root of 60=7.7

OpenStudy (jack1):

perfect and 7.75 = 2 (sqrt15) ... so ur answer should be...? ;)

OpenStudy (youngnephew34):

The square root of 15 = 3.87

OpenStudy (jack1):

yep, but just saying... ur options are A,B,C or D, so which one do u think is correct?

OpenStudy (youngnephew34):

c

OpenStudy (jack1):

The square root of 15 = 3.87 so 2 x The square root of 15 = 2 x 3.87 = 7.75 = sqrt 60

OpenStudy (jack1):

yep, perfect man!